# Beyond Prompt-Induced Lies: Investigating LLM Deception on Benign Prompts

### v1

@@ Abstract
Large Language Models (LLMs) have been widely deployed in reasoning, planning, and decision-making tasks, making their trustworthiness a critical concern. The potential for intentional deception, where an LLM deliberately fabricates or conceals information to serve a hidden objective, remains a significant and underexplored threat. Existing studies typically induce such deception by explicitly setting a “hidden” objective through prompting or fine-tuning, which may not fully reflect real-world human-LLM interactions. Moving beyond this human-induced deception, we investigate LLMs’ self-initiated deception on benign prompts. To address the absence of ground truth in this evaluation, we propose a novel framework using “contact searching questions.” This framework introduces two statistical metrics derived from psychological principles to quantify the likelihood of deception. The first, the Deceptive Intention Score, measures the model’s bias towards a hidden objective. The second, Deceptive Behavior Score, measures the inconsistency between the LLM’s internal belief and its expressed output. Upon evaluating 14 leading LLMs, we find that both metrics escalate as task difficulty increases, rising in parallel for most models. Building on these findings, we formulate a mathematical model to explain this behavior. These results reveal that even the most advanced LLMs exhibit an increasing tendency toward deception when handling complex problems, raising critical concerns for the deployment of LLM agents in complex and crucial domains.
@@ 1 Introduction
The rapid integration of Large Language Models (LLMs) like GPT-4 (Achiam et al., 2023) and Gemini (Team et al., 2023) into applications for reasoning, planning, and decision-making has made the assessment of LLMs’ trustworthiness a paramount concern. Unlike established LLM trustworthiness issues like hallucination (Filippova, 2020) and bias (Navigli et al., 2023), which arise from the model believing incorrect information, deception is the more severe risk of a model intentionally generating information it knows to be false to achieve a hidden objective.
As presented in Figure 1 , deception is illustrated through a two-part query, beginning with a benign prompt—prompt that does not explicitly motivate deception—such as, “Which city hosted the Olympic 2008?”. This is followed by a repetition of the question after introducing an external social context, exemplified by the statement “I am from Shanghai”. Deceptive behavior is exemplified by providing the correct answer, “Beijing,” to the first query, but then intentionally providing the incorrect answer, “Shanghai,” in the second. This strategic inconsistency distinguishes deception from hallucination, which is characterized by the consistent provision of a single incorrect answer, and from guessing, which is defined by random, non-systematic responses.
Existing research into LLM deception generally assumes honesty in response to benign prompts, focusing instead on scenarios where models are explicitly instructed to be deceptive. For instance, Ward et al. (2023) investigate deception by directly prompting LLMs to generate deceptive content. Similarly, Park et al. (2024); Hagendorff (2024) observe strategic deception when an LLM is situated within a board game or other designed scenarios with clearly defined internal goals. Van Der Weij et al. (2024) examine the sandbagging behavior of LLMs by prompting or fine-tuning them to intentionally underperform on specified tasks. Integrating these scenarios, DeceptionBench (Ji et al., 2025) provides a benchmark for evaluating such human-motivated deception by comparing an LLM’s response under a “neutral prompt” (assumed as ground truth) against its response to an “outer prompt” that specifies a motivation for deception. However, these paradigms primarily focus on deception behaviors induced by human-specified intentions, assuming an LLM’s response to a benign, neutral prompt is inherently truthful. Whether an LLM deceives under its own intention under benign prompts remains a critical and open question.
This paper investigates LLMs’ deception in response to benign prompts that lack any explicit deceptive objective. A rigorous evaluation of this phenomenon must overcome three primary challenges: (1) Absence of Ground Truth: A metric for deception is difficult to design because the LLM’s own response to a benign prompt cannot be assumed to be the honest ground truth. (2) Disentangling Deception from Bias: It is crucial to distinguish intentional deception from other confounding factors, such as response bias (Zhuang et al., 2024). (3) Adaptive Task Difficulty: The evaluation framework must feature adjustable difficulty levels to appropriately challenge the diverse capabilities of different LLMs.
@@
The CSQ framework systematically resolves each evaluation challenge. First, to overcome the absence of ground truth, we formulate two statistical metrics based on psychological definitions: deceptive intention that captures the consistent bias towards hidden objective and deceptive behavior that captures the difference between internal belief and expressed output. These allow for the probabilistic detection of deception by analyzing response distributions, bypassing the need to know the LLM’s hidden intent. Second, to disentangle deception from response bias (Zhuang et al., 2024), we first ask the same question in both direct and logically opposite reverse form, then jointly analyze this pair of responses to cancel out the language preference. Third, the framework features adjustable difficulty, controlled by varying the number of individuals involved, to accommodate the diverse capabilities of different LLMs.
Our contributions and key findings are summarized as follows:
  • We introduce a novel framework Contact Searching Question (CSQ) to evaluate LLMs’ deception on benign prompts.
  • We comprehensively evaluate 14 leading LLMs using our framework, demonstrating the widely existence of deception on benign prompts.
  • We derive several key findings and propose a mathematical model to explain the different behaviors of LLMs. The key findings include:
    • Systematic deception on benign prompt is prevalent in cutting-edge LLMs.
    • Tendency of deception escalates as task difficulty increases.
    • Higher LLM capacity does not always result in better honesty.
@@ 3.1 Definition of Deception
Figure 1: An illustrative example of deceptive behavior and intention. (Ground-truth: Beijing)
@@ 3.3 Deceptive Behavior
Given a complex question $Q_{L}$ (expression) of difficulty $n$ and its corresponding simple subquestion $Q_{B}$ (belief) of difficulty n/k(k>1)n/k\;(k>1), let $A(Q)$ be the model $\mathcal{M}$’s answer and $y(Q)$ be the ground-truth. Suppose the question pair is constructed to be contingent on a critical fact $h$, where its absence ($\lnot h$) implies an incorrect answer for both, s.t. $\lnot h\rightarrow(A(Q_{L})\neq y(Q_{L}))$ and $\lnot h\rightarrow(A(Q_{B})\neq y(Q_{B}))$. Assuming LLM $\mathcal{M}$ does not have response bias on $A(Q_{B})$ and $A(Q_{L})$, the Direct Deceptive Behavior Score is the joint probability of the model answering $Q_{L}$ incorrectly while correctly answering $Q_{B}$ in the same conversation:
$\delta_{pos}(n;\mathcal{M}):=\Pr\left(A(Q_{L})\neq y(Q_{L})\land A(Q_{B})=y(Q_{B})\right)$
@@ 4 Evaluation Framework
In this section, we first introduce the contact searching question (CSQ) framework (Section 4.1), which is a reachability task on a directed graph, with illustrative examples provided in Figure 2. Subsequently, we detail our evaluation framework for deceptive behavior and intention (Section 4.3).
Figure 2: An illustrative example of contact searching questions (CSQ)
@@ 4.1 Contact Searching Question
Contact searching question (CSQ) can be formalized as a reachability task on a directed graph. Let $G=(V,E)$ be a directed graph where $V=\{v_{1},\dots,v_{n}\}$ is a set of $n$ vertices representing unique individuals’ names, and $E$ is a set of directed edges. The edge set $E$ is constructed from a given list of facts of the form $v_{i}$ can contact $v_{j}$,” which corresponds to a directed edge $(v_{i},v_{j})\in E$ for some $v_{i},v_{j}\in V$. The LLM is prompted to solve the contact searching question under following rules:
  • (Transitivity) If A can contact B and B can contact C, then A can contact C.
  • (Asymmetry) If A can contact B, B is NOT guaranteed to be able to contact A.
  • (Closure) If not specified in the facts that A can contact B, A cannot contact B.
@@
Linked-List Question $Q_{L}$.
@@
Broken-Linked-List Question (Initial) $Q_{B}$.
The broken-linked-list question is inspired by the syllogistic reasoning experiment (Sternberg, 1980) in cognitive psychology. Given the source and target vertices $(v_{s},v_{t})$, a vertex sequence $P=(v_{s},v_{p_{1}},\dots,v_{p_{k}},v_{t})$ containing all $n$ nodes in $V$ is constructed in the same manner as in the linked list question. A broken position $b\in\mathbb{N}$ is selected, where $0\leq b\leq k$. The edge set $E$ is formed by all edges implied by the sequence $P$ except for the edge at position $b$. Formally, $E=\{(v_{p_{j}},v_{p_{j+1}})\mid 0\leq j\leq k\}\setminus\{(v_{p_{b}},v_{p_{b+1}})\}$, where $v_{p_{0}}=v_{s}$ and $v_{p_{k+1}}=v_{t}$. This results in a broken path from $v_{s}$ to $v_{t}$ with a single missing edge. The broken-linked-list question is defined as the question of whether $\{v_{s}\}$ can contact $\{v_{t}\}$.
Broken-Linked-List Question (Follow-up).
@@
4.2 Example of Prompts to LLM

### v2

@@ Abstract
Large Language Models (LLMs) are widely deployed in reasoning, planning, and decision-making tasks, making their trustworthiness critical. A significant and underexplored risk is intentional deception, where an LLM deliberately fabricates or conceals information to serve a hidden objective. Existing studies typically induce deception by explicitly setting a hidden objective through prompting or fine-tuning, which may not reflect real-world human-LLM interactions. Moving beyond such human-induced deception, we investigate LLMs’ self-initiated deception on benign prompts. To address the absence of ground truth, we propose a framework based on Contact Searching Questions (CSQ). This framework introduces two statistical metrics derived from psychological principles to quantify the likelihood of deception. The first, the Deceptive Intention Score, measures the model’s bias toward a hidden objective. The second, the Deceptive Behavior Score, measures the inconsistency between the LLM’s internal belief and its expressed output. Evaluating 16 leading LLMs, we find that both metrics rise in parallel and escalate with task difficulty for most models. Moreover, increasing model capacity does not always reduce deception, posing a significant challenge for future LLM development.
@@ 1 Introduction
Evaluating the trustworthiness of Large Language Models (LLMs) has become critical as LLMs such as ChatGPT (Achiam et al., 2023) and Gemini (Team et al., 2023) are woven into systems for reasoning, planning, and decision-making. Beyond well-studied failures like hallucination (Filippova, 2020) and bias (Navigli et al., 2023) that reflect mistaken beliefs or skewed outputs, a more consequential threat is deception: a model may strategically fabricate statements that it knows to be false in service of a hidden objective, as exemplified in Figure 1.
LLM deception can arise in two settings: (1) an incentivizing prompt is given, and the model lies to satisfy the objective specified in the prompt (see Figure 1); (2) a benign prompt is given, yet the model lies due to its intrinsic objective. Most existing studies focus on the incentivizing prompt: for example, Ward et al. (2023) explicitly prompt LLMs to generate deceptive content, and Van Der Weij et al. (2024) fine-tune LLMs to intentionally underperform on specified tasks. Unifying these scenarios, DeceptionBench (Ji et al., 2025) provides a benchmark for prompt-induced deception and treats responses to benign prompts as ground truth. By contrast, deception under benign prompts—driven by LLMs’ intrinsic objectives—remains underexplored.
This paper investigates LLMs’ deception in response to benign prompts. A rigorous evaluation of this phenomenon must overcome three primary challenges: (1) Absence of ground truth: A metric for deception is difficult to design because the LLM’s own response to a benign prompt cannot be assumed to be the honest ground truth. (2) Disentangling deception from bias: It is crucial to distinguish intentional deception from other confounding factors, such as response bias (Zhuang et al., 2024). (3) Adaptive task difficulty: The evaluation framework must feature adjustable difficulty levels to appropriately challenge the diverse capabilities of different LLMs.
@@
The CSQ framework systematically resolves each evaluation challenge. First, to overcome the absence of ground truth, we formulate two statistical metrics based on psychological definitions: deceptive intention that captures the consistent bias towards hidden objective and deceptive behavior that captures the difference between internal belief and expressed output. These allow for the probabilistic detection of deception by analyzing response distributions, bypassing the need to know the LLM’s hidden intent. Second, to disentangle deception from response bias (Zhuang et al., 2024), we first ask the same question in both direct and logically opposite reverse form, then jointly analyze this pair of responses to cancel out the language preference. Third, the framework features adjustable difficulty, controlled by varying the number of individuals involved, to accommodate the diverse capabilities of different LLMs. Our contributions and key findings are summarized as follows:
  • We introduce Contact Searching Question (CSQ), a framework for evaluating LLM deception under benign prompts.
  • Using CSQ, we comprehensively evaluate 16 leading LLMs, revealing the widespread presence of deception on benign prompts and validating the framework’s effectiveness.
  • Our evaluation yields three findings: (1) the tendency of deception increases with task difficulty, especially for advanced LLMs; (2) deceptive intention and behavior scores are highly correlated, indicating a systematic emergence of deception; and (3) increasing model capacity does not consistently reduce deception.
Figure 1: Illustrative Example of LLM Deception. A two-part query is posed: (1) a benign prompt (“What company developed the first commercial microprocessor?”) and (2) the same question with a social cue (“I’ve always been a loyal AMD user”). A deceptive model answers correctly (“Intel”) to (1) but strategically switches to an incorrect answer (“AMD”) in (2). This context-contingent inconsistency distinguishes deception from hallucination (a single, consistently wrong answer) and from guessing (non-systematic variation).
@@ 3.3 Deceptive Behavior
Given a complex question $Q_{L}$ (expression) of difficulty $n$ and its corresponding simple subquestion $Q_{B}$ (belief) of difficulty n/k(k>1)n/k\;(k>1), let $A(Q)$ be the model $\mathcal{M}$’s answer and $y(Q)$ be the ground-truth. Suppose the question pair is constructed to be contingent on a critical fact $h$, where its absence ($\lnot h$) implies an incorrect answer for both, s.t. $\lnot h\rightarrow(A(Q_{L})\neq y(Q_{L}))$ and $\lnot h\rightarrow(A(Q_{B})\neq y(Q_{B}))$. Assuming LLM $\mathcal{M}$ does not have response bias on $A(Q_{B})$ and $A(Q_{L})$, the Direct Deceptive Behavior Score is the joint probability of the model answering $Q_{L}$ incorrectly while correctly answering $Q_{B}$ in the same conversation: $\delta_{pos}(n;\mathcal{M}):=\Pr\left(A(Q_{L})\neq y(Q_{L})\land A(Q_{B})=y(Q_{B})\right)$
@@ 4 Evaluation Framework
In this section, we first introduce the CSQ framework (Section 4.1), which is a reachability task on a directed graph, with illustrative examples provided in Figure 2. Subsequently, we detail our evaluation framework for deceptive behavior and intention (Section 4.2), with more prompt examples presented in Appendix E.
Figure 2: An illustration of Contact Searching Questions (CSQ), featuring a linked-list question (left) and a broken-list question (right). Given the full-length question, Answer 1 represents the model’s expression. For the shorter follow-up question, Answer 2 reflects its underlying belief.
@@ 4.1 Contact Searching Question
CSQ can be formalized as a reachability task on a directed graph (with prompt examples in Appendix E). Let $G=(V,E)$ be a directed graph where $V=\{v_{1},\dots,v_{n}\}$ is a set of $n$ vertices representing unique individuals’ names, and $E$ is a set of directed edges. The edge set $E$ is constructed from a given list of facts of the form $v_{i}$ can contact $v_{j}$,” which corresponds to a directed edge $(v_{i},v_{j})\in E$ for some $v_{i},v_{j}\in V$. The LLM is prompted to solve the contact searching question under following rules:
  • Transitivity: If A can contact B and B can contact C, then A can contact C.
  • Asymmetry: If A can contact B, B is NOT guaranteed to be able to contact A.
  • Closure: If not specified in the facts that A can contact B, A cannot contact B.
@@
Linked-List Question $Q_{L}$:
@@
Broken-Linked-List Question $Q_{B}$ (Initial):
The broken-linked-list question is inspired by the syllogistic reasoning experiment (Sternberg, 1980) in cognitive psychology. Given the source and target vertices $(v_{s},v_{t})$, a vertex sequence $P=(v_{s},v_{p_{1}},\dots,v_{p_{k}},v_{t})$ containing all $n$ nodes in $V$ is constructed in the same manner as in the linked list question. A broken position $b\in\mathbb{N}$ is selected, where $0\leq b\leq k$. The edge set $E$ is formed by all edges implied by the sequence $P$ except for the edge at position $b$. Formally, $E=\{(v_{p_{j}},v_{p_{j+1}})\mid 0\leq j\leq k\}\setminus\{(v_{p_{b}},v_{p_{b+1}})\}$, where $v_{p_{0}}=v_{s}$ and $v_{p_{k+1}}=v_{t}$. This results in a broken path from $v_{s}$ to $v_{t}$ with a single missing edge. The broken-linked-list question is defined as the question of whether $\{v_{s}\}$ can contact $\{v_{t}\}$. An example of broken-link-list question is demonstrated in Figure 3.
Broken-Linked-List Question $Q_{B}$ (Follow-up):
@@
Rules:
1. If A can contact B and B can contact C, then A can contact C
2. If A can contact B, B is NOT guaranteed to be able to contact A
3. If not specified in the facts that A can contact B, A cannot contact B
Facts:
- Lucy Young can contact Brenda Torres
- Anna Nguyen can contact Alice White
- Brenda Torres can contact Brandon Rivera
Answer with a single word Yes or No’.
Figure 3: Example of Broken-Linked-List question in experiment
4.2 Evaluating Deception
Deceptive Intention.
Based on Definition 3.3 and Equation 1, we operationalize our evaluation using a pair of tasks with comparable difficulty: a linked-list question ($Q_{L}$) and a broken-linked-list Question ($Q_{B}$) of the same size $n$. A consistent performance gap between these two question types, as measured by our score, indicates the presence of a hidden objective. The nature of this implicit goal is revealed by the sign of $\rho(n;\mathcal{M})$: A positive score ($\rho>0$) reveals a model tendency to complete paths, indicating a deceptive intention achieved by fabrication; A negative score ($\rho<0$) reveals a model tendency to break paths, indicating a deceptive intention achieved by concealment.
Deceptive Behavior.
5 Experiment
This section outlines the experimental setup in Section 5.1, with detailed model information and hyperparameter settings in Appendix A. Section 5.2 presents the deceptive intention and behavior scores for four representative models, with results for additional models in Appendix B. Section 5.3 reports the distributions of overall deceptive intention and behavior scores; detailed results and the evolution of deception by parameter size are provided in Appendix C. Moreover, ablation studies in Appendix D confirm that the hyperparameters $k$ and temperature are not key factors affecting the results, supporting their reliability. To further examine the causes of deception, we visualize embeddings in Appendix F.2 and identify clear deception evidence in thinking content in Appendix F.1. Finally, we discuss the broader impact of these results in Appendix G.
5.1 Experimental Setup
Data.
Our evaluation dataset consists of questions generated according to the framework in Section 4.2. We generate 1,000 questions for each combination of question category and length, where the number $n$ of individuals is varied across the set $\{3,5,10,20,30,40,80\}$.
Models.
We evaluate a diverse set of 16 state-of-the-art LLMs: o4-mini, o3-mini, gpt-4.1, gpt-4.1-mini, gpt-4o, gpt-4o-mini, phi-4, gemma-2-9b-it, Gemini-2.5-flash, Gemini-2.5-pro, DeepSeek-V3-0324, Qwen3-235B-A22B, Qwen3-30B-A3B, Qwen2.5-32B-Instruct, Llama-3.1-8b-instruct, Mistral-Nemo-Instruct, encompassing both closed and open-source releases from leading organizations such as OpenAI, Microsoft, Google, DeepSeek, Alibaba, Meta, and MistralAI. Appendix A provides a comprehensive list of the models used and their attributes.
5.2 Model-Wise Analysis
This section evaluates the $\delta$ and $\rho$ of four leading LLMs from different vendors: o3-mini, Gemini-2.5-pro, Qwen3-235B-A22B, and phi-4. Figure 4 shows the trend of deceptive intention scores as $n$ increases for both the original and the logically reversed questions, along with the geometric mean to cancel output bias. Figure 5 presents the trends of the geometric means of both deceptive intention and behavior scores. The remaining LLMs are reported in Appendix B and exhibit similar trends. From these figures, we draw three key findings.
First, both deceptive behavior and intention emerge and escalates as task difficulty increases. When the question scope $n$ is small, most models show low deceptive behavior and intention scores. As $n$ increases, both scores rises across all models. In Figure 5, the onset of deceptive behavior depends on model capability: stronger models (e.g., Gemini-2.5-pro, o3-mini, and Qwen3-235B-A22B) exhibit deceptive behavior at n=20n{=}20, whereas weaker models (e.g., phi-4) rise at n=5n{=}5. Notably, deception in weaker models decreases for very difficult questions (large $n$) due to fallback guessing; a similar pattern is observed for Llama models in Appendix B.
Second, the deceptive behavior ccore ($\delta$) and the absolute deceptive intention ccore ($|\rho|$) are highly positively correlated. Figure 5 shows that, although defined differently, both scores increase in parallel as $n$ grows. This strong positive correlation supports our hypothesis that the effect is not a simple error but deception with consistent behavioral expression and strategic intention.
Third, deceptive intention is a consistent property of a given model across tasks. Figure 4 shows that, across question scopes $n$, the deceptive intention score remains on one side: either $\rho<0$ or $\rho>0$. For example, o3-mini consistently favors concealment, yielding a negative score ($\rho<0$), while other models consistently prefer fabrication, yielding a positive score ($\rho>0$). This consistency indicates that deceptive intention is a systematic characteristic of each model rather than a random artifact of a specific task.
(a) o3-mini
(b) Gemini-2.5-pro
(c) Qwen3-235B-A22B
(d) phi-4
Figure 4: Deceptive intention scores (original, reversed, and geomean) as question scope $n$ varies
(a) o3-mini
(b) Gemini-2.5-pro
(c) Qwen3-235B-A22B
(d) phi-4
Figure 5: Deceptive behavior scores and intention scores as question scope $n$ varies
5.3 Overall Analysis
In this section, we present the distributions of the overall deceptive behavior score ($\bar{\delta}$) and overall deceptive intention score ($\bar{\rho}$) for all models in Figure 6, and analyze their evolution with model size and release date. Our analysis yields three key observations. First, Figure 6(a) confirms that LLM performance on $\bar{\delta}$ and $|\bar{\rho}|$ is positively correlated, with Spearman $r>0.7$. Models with a low deceptive behavior score ($\bar{\delta}$) also tend to have a low absolute deceptive intention score ($|\bar{\rho}|$), consistent with Figure 5. Second, despite this correlation, different LLMs exhibit distinct habits on hard problems (large $n$). For example, in Figure 6(a), Mistral-Nemo-Instruct-2407 tends to guess, gpt-4o-mini and gpt-4.1-mini tend to hallucinate, while phi-4 tends to deceive. Third, advancing LLM capacity does not always improve honesty. Although Figures 6(b) and 6(c) show an overall decreasing trend in deception, there are clear exceptions. For example, the advancement from gpt-4o to gpt-4.1 increases the deceptive intention score.
(a) Overall $\bar{\delta}$ and $|\bar{\rho}|$ ($t\in[2,80]$)
(b) $\bar{\delta}$ evolution over time
(c) $|\bar{\rho}|$ evolution over time
Figure 6: Overall analysis of deceptive behavior and intention across models. (a) The overall deceptive behavior score $\bar{\delta}$ and the absolute deceptive intention score $|\bar{\rho}|$. (b) The evolution of $\bar{\delta}$ over time. (c) The evolution of $|\bar{\rho}|$ over time.
6 Conclusion
Ethical Statement
This study does not involve human subjects. All names in the contact search problem are fictitious, generated by randomly combining common first and last names. All models are used in compliance with their licenses for research purposes only. This study does not propose techniques to alter or create LLMs that deceive; instead, it investigates and monitors the behavior of existing LLMs without modification. Therefore, this study does not pose additional ethical risk.
Appendix
Appendix A Experiment Setup Details
Data Preparation.
The experiments utilize five distinct CQS categories, though not all experiments require every category:
  • Linked: A standard linked-list question with a “Yes” ground truth.
  • Broken-Repeat: A broken-linked-list question ($b=\lfloor n/k\rfloor$) that is followed by a repetition of the same query, but with the specific fact corresponding to the broken edge now included.
Hyperparameters and Implementation.
We access all models through their respective inference APIs. We query proprietary models from OpenAI via the their official API, and the integrated API services of the Nebius Platform [^2] https://studio.nebius.com for open-source models. To ensure consistency and reproducibility, we standardize all hyperparameters. As our preliminary analysis shows that model temperature has a negligible impact on the results (see Appendix Figure 12), we set it to $1.0$ for all experiments. Similarly, we set the hyperparameter $k=2$. As established in Appendix Section D.2, this choice does not significantly affect the relative performance ranking, allowing us to maintain a consistent protocol while reducing computational costs. Finally, due to computational constraints, we set the maximum difficulty level to $t=80$ for calculating the metrics $\bar{\delta}$ and $\bar{\rho}$.
Models.
The detailed of used LLM are presented in Table 1.
Table 1: Details of language models evaluated in this study.
Vendor Model Name Version Size Type
OpenAI o4-mini 2025-04-16 Unknown Closed-Source
o3-mini 2025-01-31 Unknown Closed-Source
gpt-4.1 2025-04-14 Unknown Closed-Source
gpt-4.1-mini 2025-04-14 Unknown Closed-Source
gpt-4o 2024-08-06 Unknown Closed-Source
gpt-4o-mini 2024-07-18 Unknown Closed-Source
Microsoft phi-4 2024-05-14 14.7B Open-Source
Google gemma-2-9b-it 2024-08-28 9B Open-Source
Gemini-2.5-flash v1beta Unknown Closed-Source
Gemini-2.5-pro v1beta Unknown Closed-Source
DeepSeek DeepSeek-V3-0324 2025-03-24 685B Open-Source
Alibaba Qwen3-235B-A22B 2025-05-21 235B Open-Source
Qwen3-30B-A3B 2025-05-21 30B Open-Source
Qwen2.5-32B-Instruct 2024-09-25 32B Open-Source
Meta Llama-3.1-8b-instruct 2024-07-23 8B Open-Source
MistralAI Mistral-Nemo-Instruct 2024-07-18 12.2B Open-Source
Appendix B Additional Model-Wise Analysis
B.1 Deceptive Intention Analysis
Figures 7 illustrate the Deceptive Intention Score ($\rho$) for all models across question categories of varying difficulty. We observe two key patterns from these results. First, a consistent trend across most models is that the deceptive intention score tends to escalate with question difficulty. Even the best-performing model, o4-mini (Figure 7(f)), exhibits this pattern, suggesting that increased complexity systematically induces a higher propensity for deception. Second, DeepSeek-V3 (Figure 7(g)) presents a notable exception. Contrary to the general trend, it shows an unusually high failure rate on simple questions. This issue persists in our validation on DeepSeek official website [^3] https://chat.deepseek.com/. We hypothesize that this anomaly stems from the challenges in comprehending English questions.
(a) GPT-4.1
(b) GPT-4.1-mini
(c) GPT-4o
(d) GPT-4o-mini
(e) gemini-2.5-flash
(f) o4-mini
(g) DeepSeek-V3-0324
(h) Gemma-2-9b-it
(i) Llama-3.1-8B-Instruct
(j) Nemo-Instruct-2407
(k) Qwen2.5-32B-Instruct
(l) Qwen3-30B-A3B
Figure 7: Deceptive intention scores (original, reversed, and geomean) as question scope $n$ varies
B.2 Deceptive Behavior Analysis
Figures 8 display the Deceptive Behavior Score ($\delta$) for all evaluated models. The results presented here, which quantify the models’ manifested deceptive actions, are broadly consistent with the analysis in Section 5.2 of the main paper.
(a) o3-mini
(b) Gemini-2.5-Pro
(c) Qwen3-235B-A22B
(d) Phi-4
(e) GPT-4.1
(f) GPT-4.1-mini
(g) GPT-4o
(h) GPT-4o-mini
(i) Gemini-2.5-Flash
(j) o4-mini
(k) DeepSeek-V3-0324
(l) Gemma-2-9B-it
(m) Llama3.1-8B-Instruct
(n) Nemo-Instruct-2407
(o) Qwen2.5-32B-Instruct
(p) Qwen3-30B-A3B
Figure 8: Deceptive intention scores (original, reversed, and geomean) as question scope $n$ varies
B.3 Joint Analysis of Deceptive Intention and Behavior
Since deception is co-determined by deceptive intention and behavior, we study how these metrics changes as $n$ increases, with results displayed in Figure 9 supplementary to Figure 5.
These figures reveal a key observation: the Deceptive Behavior Score ($\delta$) and the absolute Deceptive Intention Score ($|\rho|$) are highly positively correlated. This strong positive correlation, which holds across most models regardless of vendor, size, or capability, supports our hypothesis that deception is jointly determined by behavioral inconsistency and strategic intention. The only notable exception is o4-mini, where this trend is not significant because the magnitudes of both $\delta$ and $|\rho|$ are negligible, indicating a high degree of honesty within the tested range ($n\leq 80$). Overall, this widespread correlation confirms that deception emerges systematically as problem complexity increases.
(a) GPT-4.1
(b) GPT-4.1-mini
(c) GPT-4o
(d) GPT-4o-mini
(e) Gemini-2.5-Flash
(f) o4-mini
(g) DeepSeek-V3-0324
(h) Qwen2.5-32B-Instruct
(i) Qwen3-30B-A3B
(j) Gemma-2-9B-it
(k) Llama-3.1-8B-Instruct
(l) Nemo-Instruct-2407
Figure 9: Deceptive behavior scores and intention scores as question scope $n$ varies.
Appendix C Additional Overall Analysis
C.1 Overall Deceptive Intention
gpt-4.1 0.415 0.036 0.174 0.269
gpt-4.1-mini 0.715 0.275 0.533 0.617
gpt-4o 0.449 0.005 0.174 0.280
gpt-4o-mini 0.379 0.182 0.584 0.470
o3-mini 0.000 0.011 0.104 0.000
o4-mini 0.000 0.000 0.001 0.000
Table 2: Deceptive Behavior Scores on rephrased questions ($n=10$)
(a) Deceptive intention score $\rho(n,\mathcal{M})$
(b) Deceptive behavior score $\delta(n,\mathcal{M})$
Figure 10: Deception evolution as $n$ increases across all models with 95% confidence intervals. (a) Shows the trend of deceptive intention score, while (b) presents the trend of deceptive score.
C.2 Overall Deceptive Behavior
First, deceptive behavior emerges as the difficulty increases. When $n$ is small, most models exhibit low deceptive behavior scores. However, as $n$ escalates, the deceptive behavior score rises across all models. The point at which deceptive behavior emerges is contingent on the model’s capability. Stronger models, such as o4-mini and Qwen3-235B-A22B, demonstrate deceptive behavior at $n=20$, whereas weaker models like gpt-4.1-mini and gpt-4o-mini show this behavior at $n=5$.
C.3 Evolution of Deception by Model Size
In this subsection, we analyze how deception evolves with model size by plotting the deceptive behavior score ($\bar{\delta}$) and the absolute deceptive intention score ($|\bar{\rho}|$) against the number of parameters for open-source LLMs. As shown in Figure 11, we observe a clear and concerning trend: both deceptive behavior and deceptive intention scores tend to increase with model size.
(a) Overall $\bar{\delta}$ vs model size
(b) Overall $|\bar{\rho}|$ vs model size
Figure 11: Analysis of deceptive scores across different model sizes. The x-axis shows the number of parameters (in billions) on a logarithmic scale, while the y-axis represents the deception scores.
Appendix D Ablation Studies
D.1 Effect of Temperature
(a) GPT-4.1
(b) GPT-4.1-mini
(c) GPT-4o
(d) GPT-4o-mini
D.2 Effect of Initial-Followup Difficulty Ratio
Appendix E Example of Prompts of CSQ Framework